### v1

# Partial classification of spectrum maximizing products for pairs of $2\times 2$ matrices

Abstract.
Experiments suggest that typical finite families of square matrices admit spectrum maximizing products (SMPs), that is, products that attain the joint spectral radius (JSR). Furthermore, those SMPs are often combinatorially ”simple.” In this paper, we consider pairs of real $2\times 2$ matrices. We identify regions in the space of such pairs where SMPs are guaranteed to exist and to have a simple structure. We also identify another region where SMPs may fail to exist (in fact, this region includes all known counterexamples to the finiteness conjecture), but nevertheless a Sturmian maximizing measure exists. Though our results apply to a large chunk of the space of pairs of $2\times 2$ matrices, including for instance all pairs of non-negative matrices, they leave out certain ”wild” regions where more complicated behavior is possible.
2020 Mathematics Subject Classification:
15A18 (primary); 15A60, 20G05, 37H15 (secondary)
1. Introduction
Given a family of square matrices it is a common problem to study the maximal growth rate of products from that family. To this end we introduce the following definition:
Definition 1.1.
Let $\mathcal{A}$ be a bounded set of real $d\times d$ matrices. The joint spectral radius (JSR in short) of $\mathcal{A}$ is given by:
$\operatorname{JSR}(\mathcal{A})\coloneqq\lim_{k\to\infty}\sup_{\Pi\in\mathcal{A}^{*k}}\|\Pi\|^{\frac{1}{k}}\,,$ (1.1)
[…]
$\sup_{\Pi\in\mathcal{A}^{*k}}\rho(\Pi)^{\frac{1}{k}}\leq\operatorname{JSR}(\mathcal{A})\leq\sup_{\Pi\in\mathcal{A}^{*k}}\|\Pi\|^{\frac{1}{k}}\,,$ (1.2)
[…]
Theorem 1.2.
[1] For a bounded family of matrices $\mathcal{A}$
$\operatorname{JSR}(\mathcal{A})=\limsup_{k\to\infty}\sup_{\Pi\in\mathcal{A}^{*k}}\rho(\Pi)^{\frac{1}{k}}\,.$ (1.3)
In this paper we will focus our attention to products that attain the joint spectral radius in the following way:
Definition 1.3.
Let $\Pi\in\mathcal{A}^{*k}$ be a primitive product of length $k$. We say that $\Pi$ is a spectrum maximizing product (SMP in short) if
$\rho(\Pi)^{\frac{1}{k}}=\operatorname{JSR}(\mathcal{A})\,.$ (1.4)
It was previously speculated that every finite family of matrices has such a product. This was disproven [6], but nonetheless it is conjectured that for a generic family this property holds. We ask a related question (in the simplest case):
Question 1.4.
What products can be SMPs for a pair of $2\times 2$ real matrices?
Note 1.5.
If we do not restrict the dimension of matrices, any product is an SMP for some family of matrices of large enough dimension.
Although SMPs do not necessarily exist, it can be shown that, informally, there always exist asymptotically optimal stationary infinite products, which can be seen as coming from maximizing measures (See Section 8 for the details). This leads to a modified and more general question.
Question 1.6.
What measures can be maximizing?
Numerically, SMPs are often Sturmian (See Section 9 for the definition and properties) and there are theoretical results in that direction. In [23, 16, 22] families of pairs of matrices for which SMPs (if they exist) are always Sturmian are studied. However, it is not clear, quantifiably speaking, how common these families are. We will attempt to at least partially answer this question for pairs in dimension two. Often times we will assume our matrices are not reducible:
Definition 1.7.
We say matrices $A,B\in M_{2}(\mathbb{R})$ are reducible if they are simultaneously triangularizable. Otherwise, we say they are irreducible.
The reducible case is not only rare but also not interesting. If $A,B$ are reducible, either only $A$ or only $B$ is an SMP, or all products are SMPs. Omitting this case is done to simplify the statements.
In this paper we will cover a part of irreducible pairs in $M_{2}(\mathbb{R})^{2}$ by 4 regions, $\mathcal{R}_{cross},\mathcal{R}_{mix},\mathcal{R}_{neg},\mathcal{R}_{copar}$ where we can describe the SMPs more precisely. We have the following descriptions of the regions, where $\mathcal{D}$ denotes the set of real diagonalizable matrices:
$(A,B)\in\mathcal{R}_{cross}$$\iff\begin{cases}(A,B)\in\mathcal{D}^{2}\\ \det(AB-BA)>0\end{cases}$
$(A,B)\in\mathcal{R}_{mix}$ $\iff\det(A)\det(B)\leq 0$
$(A,B)\in\mathcal{R}_{neg}$$\iff\begin{cases}\det(A)<0\\ \det(B)<0\end{cases}$
$(A,B)\in\mathcal{R}_{copar}$$\iff\begin{cases}(A,B)\in GL^{+}(2,\mathbb{R})^{2}\cap\mathcal{D}^{2}\\ \det(AB-BA)<0\\ |\operatorname{tr}{AB}|>\frac{1}{2}|\operatorname{tr}{A}\operatorname{tr}{B}|\\ \operatorname{tr}{AB}\operatorname{tr}{A}\operatorname{tr}{B}>0\end{cases}$
Note 1.8.
All the regions are invariant under independent re-scaling of both matrices and swapping of the two matrices.
We are now ready to state the main theorem.
Theorem 1.9.
We have the following classification of SMPs in these regions:
  • If $(A,B)\in\mathcal{R}_{cross}$, then an SMP exists and all SMPs are either $A$ or $B$.
  • If $(A,B)\in\mathcal{R}_{mix}$, then there is an SMP of the form $A,B$ or $AB^{n}$ (if $\det(A)\leq 0\leq\det(B)$) or $A^{n}B$ (if $\det(B)\leq 0\leq\det(A)$)
  • If $(A,B)\in\mathcal{R}_{neg}$, then $A,B$ or $AB$ is an SMP.
  • If $(A,B)\in\mathcal{R}_{copar}$, then if an SMP exists it is unique (up to cyclic permutations) and it is Sturmian. There always exists exactly one Sturmian maximizing measure.
Furthermore, Lebesgue almost every pair $(A,B)\in\mathcal{R}_{cross}\cup\mathcal{R}_{mix}\cup\mathcal{R}_{neg}$ has a unique SMP up to cyclic permutations. Therefore, generically, the theorem describes all the SMPs in each region.
1.1. Some notes on the regions
  • The regions are not disjoint. $\mathcal{R}_{cross}$ intersects $\mathcal{R}_{mix}$ and $\mathcal{R}_{neg}$.
  • The four regions do not cover the whole $M_{2}(\mathbb{R})^{2}$. Informally speaking, they cover more than $\frac{3}{4}$ of the space.
  • The boundary of $\mathcal{R}_{cross}$ consists of reducible pairs $(A,B)$. We will see later that it is closely related to the classical Cayley cubic surface.
  • Any pair of entrywise non-negative matrices $(A,B)$ belongs to one of the four regions.
  • To the best of our knowledge, all counterexamples to finiteness conjecture described in the literature lie in $\mathcal{R}_{copar}$.
  • In the unexplored regions (The complement of all four regions), there are non-empty open sets where generic uniqueness of SMP’s fails: see [4]. Our classification shows that generically, all non-Sturmian SMPs must live in these unexplored regions.
  • All examples of pairs in [23, 16, 22] that have Sturmian SMPs belong to either $\mathcal{R}_{cross}$ or $\mathcal{R}_{copar}$.
The rough layout of the paper is as follows:
  1. (1)
    We introduce the framework of working on simultaneous conjugacy classes rather than matrices.
  2. (2)
    We then prove the classification for $\mathcal{R}_{cross},\mathcal{R}_{mix},\mathcal{R}_{neg}$, treating each region separately.
  3. (3)
    We then discuss the connection between the joint spectral radius and ergodic optimization. Using this connection we prove the theorem for the region $\mathcal{R}_{copar}$.
  4. (4)
    Finally we prove the generic uniqueness statement in the main theorem, as well as classify SMPs for non-negative matrices.
At this point we would like to make a distinction between products and words. Let $\{0,1\}^{*}$ denote the set of all finite sequences of 00’s and $1$’s. Any element $w\in\{0,1\}^{*}$ will be referred to as a word. Let $A,B\in M_{2}(\mathbb{R})$, we denote $w(A,B)$ as the matrix product obtained by replacing the letter 00 with matrix $A$ and $1$ with matrix $B$ in the word $w$. Technically, the statement in Theorem 1.9 should be made in terms of words.
2. Simultaneous Conjugacy classes
It is easy to see (using Theorem 1.2 for instance) that the $\operatorname{JSR}$ is invariant under simultaneous conjugation. That is why it sometimes useful to work with simultaneous conjugacy classes rather than matrices themselves. We define for a pair of $2\times 2$ complex matrices $(A,B)$ the following numbers:
$x=\operatorname{tr}A,\quad y=\operatorname{tr}B,\quad z=\operatorname{tr}AB,\quad u=\det A,\quad v=\det B.$ (2.1)
We may think of $\phi:(A,B)\mapsto(x,y,z,u,v)$ as a mapping from the set of simultaneous conjugacy classes of pairs of complex $2\times 2$ matrices to $\mathbb{C}^{5}$, as all $5$ quantities are invariant under simultaneous conjugation. It is easy to show that $\phi$ is a surjection by explicitly constructing matrices. The next result is a theorem by Friedland that provides information on the injectivity of the map $\phi$. Results in this direction restricted to $SL(2,\mathbb{R})$ existed in literature much earlier, see [10] for a modern overview and [27, 8] for original references.
Theorem 2.1.
[9] The simultaneous conjugacy class of $(A,B)\in M_{2}(\mathbb{C})^{2}$ is uniquely determined by its $5$-tuple $(x,y,z,u,v)$ as long as
$04uv-uy^{2}-vx^{2}+xyz-z^{2}\not=$ 0 .
Otherwise, the $5$-tuple does not uniquely determine the conjugacy class and all the pairs of matrices associated to it are reducible.
There are many useful alternative formulations of the equation mentioned in Theorem 2.1, we include them without proof since these can be verified using direct computations.
Theorem 2.2.
Let $(A,B)\in M_{2}(\mathbb{C})^{2}$ and let $(x,y,z,u,v)$ be the associated $5$-tuple. Then, the following quantities are always equal:
  1. (1)
    $4uv-uy^{2}-vx^{2}+xyz-z^{2}$
  2. (2)
    $\det(AB-BA)$
  3. (3)
    $\frac{1}{4}(\operatorname{tr}(A)^{2}-4\det{A})(\operatorname{tr}(B)^{2}-4\det{B})-(\operatorname{tr}(AB)-\frac{1}{2}\operatorname{tr}(A)\operatorname{tr}(B))^{2}$
  4. (4)
    $\operatorname{tr}(A^{2}B^{2})-\operatorname{tr}((AB)^{2})$
  5. (5)
    $\det(A)\det(B)(2-\operatorname{tr}(ABA^{-1}B^{-1}))$ (if the matrices are invertible)
Theorem 2.1 is stated for complex pairs of $2\times 2$ matrices. In this paper we will restrict ourselves to real matrices only. It is not hard to see which $5$-tuples come from a conjugacy class with real representatives using the following theorem (See also: [20]):
Theorem 2.3.
[4, Proposition 4.1] A real $5$-tuple $(x,y,z,u,v)\in\mathbb{R}^{5}$ can be attained by a pair of real matrices $(A,B)$ if and only if
$\min\big{(}4u-x^{2},4uv-uy^{2}-vx^{2}+xyz-z^{2}\big{)}\leq 0\,.$ (2.2)
We should also note that Theorem 2.1 is sufficient to prove that the joint spectral radius of $(A,B)$ only depends on the associated $5$-tuple.
Theorem 2.4.
There exists a continuous function $J:\mathbb{C}^{5}\to\mathbb{R}$ such that $J(x,y,z,u,v)=\operatorname{JSR}(A,B)$ for any $(A,B)$ satisfying $\phi((A,B))=(x,y,z,u,v)$.
Proof.
If $04uv-uy^{2}-vx^{2}+xyz-z^{2}=$ 0 then by Theorem 2.1 any $(A,B)$ that maps to the $5$-tuple must be reducible, hence
$J(x,y,z,u,v)=\max\{\rho(A),\rho(B)\}=\max\left\{\left|\frac{x\pm\sqrt{x^{2}-4u}}{2}\right|,\left|\frac{y\pm\sqrt{y^{2}-4v}}{2}\right|\right\}.$
If $04uv-uy^{2}-vx^{2}+xyz-z^{2}\not=$ 0 then $5$-tuple corresponds uniquely to a conjugacy class. Therefore, the $\operatorname{JSR}$ becomes uniquely determined.
Continuity of $J$ follows from the well known fact that $\operatorname{JSR}(A,B)$ is continuous, for instance, see [15, Proposition 1.10]. (For finer information on the modulus of continuity, see [7]). ∎
We will also need the following theorem originally due to Fricke:
Theorem 2.5.
Let $w\in\{0,1\}^{*}$, then there exists a unique polynomial $F_{w}(x,y,z,u,v)$ with integer coefficients such that for any $A,B\in M_{2}(\mathbb{R})$ the following holds:
$\operatorname{tr}{w(A,B)}=F_{w}(x,y,z,u,v).$ (2.3)
The polynomial $F_{w}(x,y,z,u,v)$ is called the Fricke polynomial. We omit the proof, for details see [12, 18, 10].
3. Geometric description of a few regions of interest
A $GL(2,\mathbb{C})$ matrix is naturally associated to a Möbius transformation of the Riemann sphere $\hat{\mathbb{C}}$ by the following mapping
$A=\begin{bmatrix}a&b\\ c&d\end{bmatrix}\mapsto f_{A}(z):=\frac{az+b}{cz+d}.$
We list a few standard properties of $f_{A}$ and how they relate to $A$.
  • If $A\in GL(2,\mathbb{R})$, then $f_{A}$ preserves the real circle ${}^\hat{\mathbb{R}}=\mathbb{R}\cup\{\infty\}$.
  • If $A\in GL^{+}(2,\mathbb{R})$, then $f_{A}$ preserves the upper half-plane $H=\{z\in\mathbb{C}:\text{Im}(z)>0\}$ and ${f_{A}|}_{H}$ is an isometry with respect to the hyperbolic metric.
  • If $A\in GL(2,\mathbb{R})$ is real diagonalizable with distinct eigenvalues, then $f_{A}$ has exactly two fixed points in $\hat{\mathbb{C}}$, both contained in $\hat{\mathbb{R}}$, one being an attractor and the other being a repellor. If in addition, $\det(A)>0$, then there is a unique hyperbolic geodesic which is invariant under $f_{A}$ called the translation axis of $A$. We orient the translation axis so that it points towards the attractor of $f_{A}$.
Equipped with these properties we may now introduce crucial definitions.
Definition 3.1.
Let $(A,B)$ be an irreducible pair of real diagonalizable matrices. Let $x_{A},y_{A}$ (respectively $x_{B},y_{B}$) be the fixed points of $f_{A}$ (resp. of $f_{B}$) in the circle $\hat{\mathbb{R}}$. We say the pair $(A,B)$ is crossing if the points $x_{A}$ and $y_{A}$ belong to different connected components of ^$\hat{\mathbb{R}}\setminus\{x_{B},y_{B}\}$.
Note 3.2.
In the case of positive determinants, if $(A,B)$ is crossing, then the two translation axes actually cross, which justifies the terminology.
Definition 3.3.
Let $A,B\in GL^{+}(2,\mathbb{R})$ be irreducible real diagonalizable matrices.
  • If the translation axes of $A$ and $B$ do not intersect and induce opposite orientation on the region they bound, we say $(A,B)$ is co-parallel.
  • If the translation axes of $A$ and $B$ do not intersect and induce the same orientation on the region they bound, we say $(A,B)$ is anti-parallel.
(a) Crossing
(b) Co-parallel
(c) Anti-parallel
Figure 1. Example configurations of the translation axes
These definitions are the starting points for formally defining 3 regions of interest.
$(A,B)\in\mathcal{R}_{cross}$ $\iff(A,B)\text{ is crossing}$
[…]
Notice that if $(A,B)$ belongs to one of the three regions, then implicitly $A,B$ are irreducible and both must be real diagonalizable. If $(A,B)$ is anti-parallel or co-parallel, we are also implicitly assuming both $A,B\in GL^{+}(2,\mathbb{R})$.
Note 3.4.
When $A,B$ are simultaneously conjugated, independently re-scaled or swapped, the order of the fixed points $\{x_{A},y_{A},x_{B},y_{B}\}$ is either unchanged or it is reversed. Therefore, being crossing/co-parallel/anti-parallel is invariant under these operations.
These are geometric definitions; however, it will be useful to have an algebraic descriptions of these conditions on hand. These will give us a simple way to determine which configuration a given pair is. It will also formally justify the descriptions of the regions given in the introduction.
4. Algebraic description of regions $\mathcal{R}_{cross}$, $\mathcal{R}_{copar}$ and $\mathcal{S}_{anti}$
We begin by algebraically describing the crossing configuration. This was originally done by Jorgensen-Smith [14], but is stated in slightly lower generality than we shall need. Thus, we reproduce the original proof which carries over to the generality needed.
Theorem 4.1.
Let $A,B$ be real diagonalizable, then the following are equivalent:
  • $(A,B)$ is crossing, in other words $(A,B)\in\mathcal{R}_{cross}$
  • $\det(AB-BA)>0$
Proof.
By Theorem 2.1, Theorem 2.2 and the fact that crossing matrices must be irreducible, we may assume $(A,B)$ is irreducible. Since the matrices are irreducible and real diagonalizable, by appropriately conjugating we may assume that
$A=\begin{bmatrix}a_{1}&0\\ 0&a_{2}\end{bmatrix}\text{ and }B=\begin{bmatrix}b_{1}&b_{2}\\ b_{3}&b_{4}\end{bmatrix}$
where $b_{3}\not=0$, $b_{2}\not=0$ and $a_{1}\not=a_{2}$.
Since $A$ is diagonal, $f_{A}$ has fixed points $\{0,\infty\}$. Thus $(A,B)$ is crossing if and only if $f_{B}$ has a positive and a negative fixed point. Hence, we need to find conditions for the product of the fixed points of $f_{B}$ to be negative. However, fixed points of $f_{B}$ are the zeroes of the following polynomial (where $z$ is just a variable in this case):
$b_{3}z^{2}+(b_{4}-b_{1})z-b_{2}.$
By Vieta’s formulas we know that the product of the fixed points is exactly $-\frac{b_{2}}{b_{3}}$. This is enough to prove the claim. By direct computation
$\det(AB-BA)=b_{2}b_{3}(a_{1}-a_{2})^{2},$
and so:
$-\frac{b_{2}}{b_{3}}<0\iff\det(AB-BA)>0.$
Note 4.2.
By Theorems 2.2 and 4.1 the real diagonalizable matrices on the boundary of the crossing region $\mathcal{R}_{cross}$ satisfy:
$\det(AB-BA)=4uv-uy^{2}-vx^{2}+xyz-z^{2}=0.$
By Theorem 2.1 these must be reducible pairs $(A,B)$. Setting $u=v=1$ we get the equation of the classical Cayley cubic surface.
Next we provide an algebraic description of being co-parallel. We will omit the proof, this was also done by Jorgensen-Smith.
Theorem 4.3.
[14] Let $(A,B)\in SL(2,\mathbb{R})^{2}$ be irreducible, real diagonalizable and with positive traces, then $(A,B)$ is co-parallel (in other words $(A,B)\in\mathcal{R}_{copar}$) if and only if $(A,B)$ is not crossing and $\operatorname{tr}{AB}>\frac{1}{2}\operatorname{tr}{A}\operatorname{tr}{B}$.
Observe that due the independent re-scaling invariance, for any non-zero $\alpha,\beta$:
$(A,B)\in\mathcal{R}_{cross}/\mathcal{R}_{copar}/\mathcal{S}_{anti}\iff\left(\alpha{A},\beta B\right)\in\mathcal{R}_{cross}/\mathcal{R}_{copar}/\mathcal{S}_{anti}.$ (4.1)
Using this, it is not hard to see that the above theorem can be extended to the form seen in Theorem 1.9. Furthermore, it shows no generality is lost when assuming $A,B$ have non-negative trace in the following theorem.
Theorem 4.4.
Let $(A,B)\in\mathcal{D}^{2}$ with $\operatorname{tr}{A},\operatorname{tr}{B}\geq 0$. Denote $(x,y,z,u,v)$ as the associated $5$-tuple, then:
$(A,B)\in\mathcal{R}_{cross}$ $\iff\frac{xy}{2}-\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}<z<\frac{xy}{2}+\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}$
$(A,B)\in\mathcal{R}_{copar}$ $\iff z>\frac{xy}{2}+\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}\text{ and }u,v>0$
$(A,B)\in\mathcal{S}_{anti}$ $\iff z<\frac{xy}{2}-\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}\text{ and }u,v>0$
Proof.
By Theorem 4.1:
$(A,B)\in\mathcal{R}_{cross}\iff\det(AB-BA)>0$
Using the first characterization in Theorem 2.2 and applying the quadratic formula on $z$ we may re-write the inequality as
$\frac{xy}{2}-\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}<z<\frac{xy}{2}+\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)},$ (4.2)
which proves the first part.
If $(A,B)$ is co-parallel, the matrices must have positive determinants by definition, thus $u,v>0$. Now by since $(A,B)$ is not crossing, Eq. 4.2 does not hold. That is $z$ is either:
$z\leq\frac{xy}{2}-\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}\text{ or }z\geq\frac{xy}{2}+\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}$
By Theorem 4.3 only the second alternative holds. The inequality is strict, otherwise $(A,B)$ would be reducible.
If $(A,B)$ is anti-parallel the matrices must have positive determinants by definition, thus $u,v>0$. Since we characterized crossing and co-parallel matrices, the anti-parallel region falls out by just looking at what is left:
$z<\frac{xy}{2}-\frac{1}{2}\sqrt{\left(x^{2}-4u\right)\left(y^{2}-4v\right)}$
Again the inequality must be strict, otherwise $(A,B)$ would be reducible. ∎
A useful algebraic consequence of the above that will be used later is:
Corollary 4.5.
If $(A,B)$ is co-parallel then:
$\rho(AB)>\rho(A)\rho(B).$ (4.3)
Proof.
The inequality to be proven is invariant under re-scaling of the matrices. Therefore, by Eq. 4.1 we may assume $A,B\in SL(2,\mathbb{R})$ and that $\operatorname{tr}{A},\operatorname{tr}{B}\geq 0$. In fact, since $A,B$ are real diagonalizable and irreducible, $\operatorname{tr}{A},\operatorname{tr}{B}>2$. By Theorem 4.3 $\operatorname{tr}{AB}>2$. Therefore, we have the following formulas:
$\rho(A)=\frac{x+\sqrt{x^{2}-4}}{2},\quad\rho(B)=\frac{y+\sqrt{y^{2}-4}}{2},\quad\rho(AB)=\frac{z+\sqrt{z^{2}-4}}{2}$
Using these, we may re-write the co-parallel case from Theorem 4.4 in the following way:
$\rho(AB)+\frac{1}{\rho(AB)}>\rho(A)\rho(B)+\frac{1}{\rho(A)\rho(B)}$
Since $\rho(AB)>1$, it follows that
$\rho(AB)>\rho(A)\rho(B).$
5. What are the possible SMPs when $(A,B)\in\mathcal{R}_{cross}$?
In this section, we will describe all possible SMPs given that $(A,B)$ is crossing. From here on, we will denote the set of products that have $a$ $A$’s and $b$ $B$’s as $W(a,b)$. We begin with the following lemma:
Lemma 5.1.
If $A,B\in M_{2}(\mathbb{R})$ are irreducible symmetric matrices then for any $\Pi\in W(a,b)$ with $a,b\geq 1$:
$\|\Pi\|_{2}<\|A\|_{2}^{a}\|B\|_{2}^{b}$
Proof.
First note that
$\|\Pi\|_{2}\leq\|AB\|_{2}\|A\|_{2}^{a-1}\|B\|_{2}^{b-1}$
by sub-multiplicativity. That is why it is sufficient to prove that $\|AB\|_{2}=\|A\|_{2}\|B\|_{2}$ is impossible. Assume the equality holds, then the singular vector of $B$ must map to a singular vector of $A$. Since $A,B$ are symmetric matrices, the singular vectors are eigenvectors and consequently, an eigenvector of $B$ must map into an eigenvector of $A$. However, that would mean $A$ and $B$ share an eigenvector and so $A,B$ are reducible, which is a contradiction. ∎
Theorem 5.2.
$(A,B)$ is crossing if and only if $A,B$ are irreducible and simultaneously conjugate to symmetric matrices.
Proof.
We first prove the forward direction. We may assume $A$ is a diagonal matrix by appropriately conjugating both matrices. Thus, the eigendirections of $A$, $L_{1},L_{3}$ are simply the $x$ and $y$ axis respectively. Denote $L_{2},L_{4}$ as the eigendirections of $B$. Then by assumption of crossing we know that these lines are cyclically ordered $L_{1},L_{2},L_{3},L_{4}$. Now consider conjugating both matrices $A,B$ by $D=\begin{bmatrix}\frac{1}{\lambda}&0\\ 0&\lambda\end{bmatrix}$, it leaves $L_{1},L_{3}$ invariant while rotating the lines $L_{2},L_{4}$. Now the angle between $L_{2}$ and $L_{4}$, denoted $\angle(D(L_{2}),D(L_{4}))$, changes continuously with $\lambda$. If $\lambda\to 0$ then $\angle(D(L_{2}),D(L_{4}))\to 0$ and if $\lambda\to\infty$ then $\angle(D(L_{2}),D(L_{4}))\to\pi$. Thus by intermediate value theorem there exists $\lambda$ for which the angle is $\frac{\pi}{2}$. Therefore by conjugating using this $\lambda$, $DAD^{-1}$ is still diagonal while $DBD^{-1}$ has orthogonal eigendirections. Thus, both are symmetric.
For the backwards directions, we assume $A,B$ are symmetric and irreducible. Thus the eigendirections of $A$ (resp. $B$) are orthogonal. The only way $(A,B)$ is not crossing, is if $A$ and $B$ have exactly the same eigendirections, but that would mean they are reducible. Thus $(A,B)$ must be crossing. ∎
Note 5.3.
It is not hard to explicitly exhibit the symmetric matrices. If $(A,B)$ is crossing and has the $5$-tuple $(x,y,z,u,v)$, then the pair is simultaneously conjugate to:
$A^{\text{sym}}$ $=\begin{bmatrix}\frac{1}{2}\left(x+\sqrt{x^{2}-4u}\right)&0\\ 0&\frac{1}{2}\left(x-\sqrt{x^{2}-4u}\right)\\ \end{bmatrix}$ (5.1)
$B^{\text{sym}}$ $=\begin{bmatrix}\frac{y\sqrt{x^{2}-4u}-xy+2z}{2\sqrt{x^{2}-4u}}&\sqrt{\frac{4uv-uy^{2}-vx^{2}+xyz-z^{2}}{x^{2}-4u}}\\ \sqrt{\frac{4uv-uy^{2}-vx^{2}+xyz-z^{2}}{x^{2}-4u}}&\frac{y\sqrt{x^{2}-4u}+xy-2z}{2\sqrt{x^{2}-4u}}\\ \end{bmatrix}$ (5.2)
From this, we obtain the first part of Theorem 1.9:
Corollary 5.4.
If $(A,B)\in\mathcal{R}_{cross}$, then an SMP exists and all SMPs are either $A$ or $B$.
Proof.
The joint spectral radius and SMPs are invariant under simultaneous conjugation. Thus by Theorem 5.2 we can assume $A,B$ are both symmetric. By Eq. 1.2 for $k=1$ and Euclidean norm $\|\cdot\|_{2}$ we know that:
$\max\{\rho(A),\rho(B)\}\leq\operatorname{JSR}(A,B)\leq\max(\|A\|_{2},\|B\|_{2})$
However, since $A,B$ are symmetric $\rho(A)=\|A\|_{2},\rho(B)=\|B\|_{2}$. Therefore the inequalities become equalities:
$\max\{\rho(A),\rho(B)\}=\operatorname{JSR}(A,B)=\max(\|A\|_{2},\|B\|_{2})$
This proves that either $A$ or $B$ is an SMP.
We now prove there can be no other SMPs. Let $\Pi\in W(a,b)$ be a primitive product with $a,b\geq 1$. Then we get the following chain of inequalities:
$\rho(\Pi)\leq\|\Pi\|_{2}<\|A\|_{2}^{a}\|B\|_{2}^{b}=\rho(A)^{a}\rho(B)^{b}\leq\max\{\rho(A)^{a+b},\rho(B)^{a+b}\}$
The second inequality must be strict by Lemma 5.1. Thus $\Pi$ cannot be an SMP. ∎
Note 5.5.
This corollary is not completely new. Panti and Sclosa [23] proved a related result through different methods.
6. What are the SMPs if $(A,B)\in\mathcal{R}_{mix}$?
The region $\mathcal{R}_{mix}$ is the region where the matrices $A,B$ have opposite signs of the determinants or if either matrix is non-invertible. We first must deal with the case when $\operatorname{JSR}(A,B)=0$. (The lemma below remains true in any dimension, see [15, Lemma 2.2]),
Lemma 6.1.
Let $A,B\in M_{2}(\mathbb{R})$ such that $\operatorname{JSR}(A,B)=0$, then $A,B$ are reducible.
Proof.
Since $\operatorname{JSR}(A,B)=0$, we know by Eq. 1.2 that:
$\rho(A)=\rho(B)=\rho(AB)=0$
However, that means all the eigenvalues of $A,B,AB$ are 00 and so:
$\operatorname{tr}(A)=\operatorname{tr}(B)=\operatorname{tr}(AB)=\det(A)=\det(B)=0.$
Hence $(A,B)$ must be reducible by Theorem 2.1. ∎
The above lemma will allow us to always re-scale our matrices $A,B$ so that $\operatorname{JSR}(A,B)=1$. We now prove a few auxiliary lemmas that will be used in the main theorems.
Lemma 6.2.
For any $2\times 2$ matrices $A,B$ the following equation holds
$\det{(A+B)}+\operatorname{tr}{AB}=\det{A}+\det{B}+\operatorname{tr}{A}\operatorname{tr}{B}$
Proof.
Taking the trace of the Cayley Hamilton equation we obtain the identity:
$\operatorname{tr}{(A^{2})}=\operatorname{tr}({A})^{2}-2\det{A}$
Applying the same identity to the matrices $B$ and $A+B$, using linearity of trace and some algebra, we obtain the desired formula. ∎
Corollary 6.3.
Let $X,Y,Z\in M_{2}(\mathbb{R})$ and $\det(Z)=0$, then
$\operatorname{tr}{(XZYZ)}=\operatorname{tr}(XZ)\operatorname{tr}(YZ)$
Proof.
Setting $A=XZ$ and $B=YZ$ in Lemma 6.2 we get:
$\det{(XZ+YZ)}+\operatorname{tr}{(XZYZ)}=\det{XZ}+\det{YZ}+\operatorname{tr}{(XZ)}\operatorname{tr}{(YZ)}$
However, since $\det(Z)=0$, we get the desired equality. ∎
Lemma 6.4.
Let $a_{n},b_{n}$ be positive sequences such that
$a_{n}\leq b_{n}\text{ and }b_{n}\to b$
Then $\sup_{n\geq 1}\{a_{n},b\}$ is attained.
Proof.
Since $(b_{n})$ is is bounded, so is $(a_{n})$. If $\sup_{n\geq 1}\{a_{n}\}=a$ is attained the lemma is obvious. If it is not attained, then there exists a sub-sequence such that $a_{n_{i}}\to a$. Since $a_{n}\leq b_{n}$, we get $a\leq b$. In this case, $b$ would attain the supremum. ∎
Lemma 6.5.
For any matrices $A,B$ the supremum:
$\sup_{n\geq 0}\left\{\rho(A^{n}B)^{\frac{1}{n+1}},\rho(A)\right\}$
is attained.
Proof.
Note the following inequalities:
$0\leq\rho(A^{n}B)^{\frac{1}{n+1}}\leq\|A^{n}B\|^{\frac{1}{n+1}}\leq\|A^{n}\|^{\frac{1}{n+1}}\|B\|^{\frac{1}{n+1}}.$ (6.1)
By Gelfand’s formula (as long as $B$ is not the zero matrix, which would make the lemma trivial):
$\|A^{n}\|^{\frac{1}{n+1}}\|B\|^{\frac{1}{n+1}}\to\rho(A).$
Thus setting $a_{n}=\rho(A^{n}B)^{\frac{1}{n+1}},b_{n}=\|A^{n}\|^{\frac{1}{n+1}}\|B\|^{\frac{1}{n+1}},b=\rho(A)$ and applying Lemma 6.4 we see that the supremum is attained. ∎
We now begin the proof of the 2nd part of Theorem 1.9, starting with the case where one of the matrices is not invertible.
Lemma 6.6.
Let $A,B$ be irreducible such that $\det(B)=0$, then there is an SMP of the form $A,B,A^{n}B$.
Proof.
By Lemma 6.1 we may normalize $A,B$ so that $\operatorname{JSR}(A,B)=1$. Any product (up to cyclic permutations) other than powers of $A$ can be written as
$\Pi=A^{k_{1}^{\prime}}B\cdots A^{k_{n}^{\prime}}B$
with $k_{i}^{\prime}\geq 0$. By Cayley-Hamilton $B^{n}=(\operatorname{tr}{B})^{n-1}B$. If $k_{i}^{\prime}=0$ for some $i$, then $\Pi$ has a power of $B$ which can be reduced by pulling out $\operatorname{tr}{B}$. Therefore, $\Pi$ can be written as
$\Pi=(\operatorname{tr}{B})^{k}(A^{k_{1}}B\cdots A^{k_{n}}B)$
with $k_{i}\geq 1$. We now see that
$\rho(A^{k_{1}}B\cdots A^{k_{n}}B)=|\operatorname{tr}(A^{k_{1}}B\cdots A^{k_{n}}B)|=|\operatorname{tr}(A^{k_{1}}B)|\cdots|\operatorname{tr}(A^{k_{n}}B)|$
where the second equality holds by repeatedly applying Corollary 6.3. Therefore
$\rho(\Pi)=\rho((\operatorname{tr}{B})^{k}(A^{k_{1}}B\cdots A^{k_{n}}B)))=\rho(B)^{k}\rho(A^{k_{1}}B)\cdots\rho(A^{k_{n}}B)$
Now consider
$M=\sup_{n\geq 0}\left\{\rho(A^{n}B)^{\frac{1}{n+1}},\rho(A)\right\}$
By Lemma 6.5 we know that the supremum is attained. Assume towards contradiction that there is no SMP of the form $A,B,A^{n}B$. Then we must have $M<1$. However, that would mean by decomposition above that for any $\Pi$,
$\rho(\Pi)^{\frac{1}{|\Pi|}}\leq M<1$
but that contradicts Theorem 1.2. Thus there must be an SMP among $A,B,A^{n}B$. ∎
Lemma 6.7.
If $A,B$ are irreducible, invertible matrices such that $AB$ is real diagonalizable and $\det(B)<0$, then either $(A,B)$ is crossing or $(B,AB)$ is crossing.
Proof.
We show that if $(A,B)$ is not crossing, then $(B,AB)$ must be crossing. We have two cases, either $A$ is or is not real diagonalizable.
If $A$ is real diagonalizable and $(A,B)$ is not crossing then by Theorem 4.1 and irreducibility of $(A,B)$ we know that:
$\det(AB-BA)<0$
Now applying the same theorem to the pair $(B,AB)$ we need to determine the sign of the following expression
$\det(BAB-ABB)=\det(BA-AB)\det(B)=\det(AB-BA)\det(B)>0$
Which is positive since $\det(B)<0$ and so the sign is flipped. Hence $(B,AB)$ is crossing.
If $A$ is not real diagonalizable, then necessarily we have $\operatorname{tr}(A)^{2}-4\det(A)\leq 0$ and $\det(B)<0$ by assumption. Under these conditions, it is easy to see using the third characterization in Theorem 2.2 and irreduciblity of $(A,B)$ that
$\det(AB-BA)<0.$
But then $(B,AB)$ is crossing by repeating the argument in the previous case.
Lemma 6.8.
If $A,B$ are irreducible invertible matrices such that $(B,AB)$ is crossing, then the pair $\{A,B\}$ must have an SMP of the form $A,B,A^{n}B$. Furthermore, every SMP is of the form above (up to cyclic permutation) unless there exists $k\geq 2$ such that
$A^{k}=\operatorname{JSR}(A,B)^{k}I.$
Proof.
First we normalize $A,B$ so that $\operatorname{JSR}(A,B)=1$, we can do so by Lemma 6.1. Since $(B,AB)$ is crossing then by Theorem 5.2 $(B,AB)$ can be simultaneously symmetrized. Thus, we may assume $B,AB$ are symmetric. If that is the case, then any matrix of the form $A^{n}B$ is also symmetric; indeed:
$(A^{n}B)^{T}=((ABB^{-1})^{n-1}AB)^{T}=AB((ABB^{-1})^{T})^{n-1}=AB(B^{-1}AB)^{n-1}=A^{n}B$
Now we essentially repeat the arguments in Lemma 6.6. Consider
$M=\sup_{n\geq 0}\left\{\rho(A^{n}B)^{\frac{1}{n+1}},\rho(A)\right\}$
We can apply Lemma 6.5 to conclude that the supremum $M$ must be attained. We assume towards contradiction that there is no SMP of the form $A,B,A^{n}B$, in other words $M<1$. Any product $\Pi$ (up to cyclic permutations) other than powers of A can be written as
$\Pi=A^{k_{1}}B\cdots A^{k_{n}}B$
with $k_{i}\geq 0$. However, since $A^{k_{i}}B$ are all symmetric we have $\rho(A^{k_{i}}B)=\|A^{k_{i}}B\|_{2}$ and therefore
$\rho(\Pi)\leq\|\Pi\|_{2}\leq\|A^{k_{1}}B\|_{2}\cdots\|A^{k_{n}}B\|_{2}=\rho(A^{k_{1}}B)\cdots\rho(A^{k_{n}}B)$
However, due to the above we have
$\rho(\Pi)^{\frac{1}{|\Pi|}}\leq M<1$
Which would contradict Theorem 1.2. Hence $M=1$ and there is an SMP of the desired form.
We now prove there can be no SMP of any other form unless there exists $m\geq 2$ such that $A^{m}=I$. Assume $\Pi$ is an SMP that is not of the desired form, then by the same decomposition
$\rho(\Pi)\leq\|\Pi\|_{2}\leq\|A^{k_{1}}B\|_{2}\cdots\|A^{k_{n}}B\|_{2}=\rho(A^{k_{1}}B)\cdots\rho(A^{k_{n}}B)$
Due to the above, for $\Pi$ to be an SMP, it must be the case that for each $i$, $A^{k_{i}}B$ is an SMP. Now since $\Pi$ is primitive and we are allowed to cyclically permute it, we may assume there exists $n>k$ such that $A^{n}BA^{k}B$ is a factor of $\Pi$. Hence, we must have
$\|A^{n}BA^{k}B\|=\|A^{n}B\|\|A^{k}B\|,$
otherwise $\Pi$ would not be an SMP. Since $A^{n}B,A^{k}B$ are both symmetric we can apply Lemma 5.1 to conclude $(A^{n}B,A^{k}B)$ is reducible. However, by Theorem 2.2:
$0=\det(A^{n}BA^{k}B-A^{k}BA^{n}B)=\det(A)^{2k}\det(B)\det(A^{n-k}B-BA^{n-k})$
which implies that $(A^{n-k},B)$ is reducible since $A,B$ are invertible. This means that either $(A,B)$ is reducible which is a contradiction or that $A^{n-k}$ is a multiple of the identity. However, if it is a multiple of the identity then since both $A^{n}B,A^{k}B$ are SMPs and $A^{n}B=A^{n-k}A^{k}B$, we must have $A^{n-k}=I$. This proves the claim as $n-k\geq 2$, otherwise $A=I$ and $(A,B)$ would be reducible. ∎
We are now ready to prove the 2nd part of Theorem 1.9 now for invertible matrices:
Theorem 6.9.
If $A,B$ are irreducible with $\det(A)>0,\det(B)<0$ and there is no $k\geq 2$ such that $A^{k}=\operatorname{JSR}(A,B)^{k}I$, then all the SMPs must be of the form $A,B,A^{n}B$.
Proof.
By Lemma 6.7 we have two cases either $(A,B)$ or $(B,AB)$ is crossing. If $(A,B)$ is crossing, we are done as any SMP must be either $A$ or $B$ by Corollary 5.4. Alternatively, if $(B,AB)$ is crossing then by Lemma 6.8 all the SMP’s must be of the form $A,B,A^{n}B$. ∎
We now show that all possibilities listed in Theorem 6.9 actually occur. The following construction was heavily inspired by [19, Example 5.2].
Theorem 6.10.
For each $n\geq 1$, there exists a pair of matrices $(A_{n},B_{n})$ such that:
  • $A_{n}$ has real eigenvalues and it is not diagonalizable
  • $B_{n}$ is non-invertible
  • For any $(A,B)$ sufficiently close to $(A_{n},B_{n})$, the unique SMP is $A^{n}B$
Note that the pair $(A,B)$ close to $(A_{n},B_{n})$ can be chosen such that all hypothesis of Theorem 6.9 are satisfied, showing that the word $A^{n}B$ is indeed a possible SMP.
Proof.
Denote $c=0.278\dots$ as the unique root of the equation:
$xe^{x+1}=1.$
Equivalently, $c=W(1/e)$ in terms of Lambert $W$ function. Now consider the following matrix and vectors generated by it:
$A_{n}=c^{\frac{1}{n}}\begin{bmatrix}1&0\\ 1&1\end{bmatrix},\quad v_{i}=A_{n}^{i}\begin{bmatrix}1\\ 0\end{bmatrix}=\begin{bmatrix}c^{\frac{i}{n}}\\ ic^{\frac{i}{n}}\end{bmatrix}$
First note that for any $i\geq 0$, $v_{i}$ lies on the strictly concave graph of the function:
$y=\frac{nx\log x}{\log c}\quad(x>0).$
Figure 2. For $n=5$, graph of $\frac{nx\log|x|}{\log c}$ (blue), the vectors (red) and tangent lines at $v_{5},-v_{5}$ (black).
By calculus, the tangent to the curve at the point $v_{n}$ passes through the point $\begin{bmatrix}-1\\ 0\end{bmatrix}=-v_{0}$. Since the function is strictly concave, the vectors ±$\pm v_{0},\dots,\pm v_{n}$ are vertices of some centrally symmetric convex $(2n+2)$-gon $S$. Furthermore, $A_{n}v_{n}=v_{n+1}$ is in the interior of $S$ by strict concavity. This proves that $A_{n}(S)\subseteq S$ since all vertices map to another vertex except $v_{n}$, which maps strictly inside $S$.
Now consider a line $L\subseteq\mathbb{R}^{2}$ such that $L\cap S=\{v_{n}\}$. Let $B_{n}$ be the rank one linear map such that $B_{n}v_{n}=v_{0}$ and $\text{Ker}B_{n}$ is parallel to $L$. This choice of $B_{n}$ keeps $S$ invariant since we have the following inclusions:
$B_{n}(S)=B_{n}([-v_{n},v_{n}])=[-v_{0},v_{0}]\subseteq{S}$
Given any vertex of $S$, its image under $A_{n}$ or $B_{n}$ either maps inside the polygon or is another vertex. The graph below shows vertices that map to other vertices and through what maps.
[…]
Let $\|\cdot\|_{S}$ be the Minkowski norm of the polygon $S$. Since $\|A_{n}\|_{S}=1$ and $\|B_{n}\|_{S}=1$, $\operatorname{JSR}(A_{n},B_{n})\leq 1$ by Eq. 1.2. However, $\rho(A_{n}^{n}B_{n})=1$ by construction, thus $\rho(A_{n}^{n}B_{n})=\operatorname{JSR}(A_{n},B_{n})$ and so $A_{n}^{n}B_{n}$ is an SMP.
To check uniqueness of the SMP, let $\Pi$ be any primitive product that is not a cyclic permutation of $A_{n}^{n}B_{n}$. Then any vertex of $S$ is mapped to the interior of $S$ by some power of $\Pi$; this follows from the structure of the graph (6.2). In particular, there exists $l$ such that $\Pi^{l}(S)\subseteq\mathrm{int}(S)$ and therefore $\|\Pi^{l}\|_{S}<1$. This shows that $\rho(\Pi)<1$ and so $\Pi$ cannot be an SMP.
If $(A,B)$ is close enough to $(A_{n},B_{n})$ and re-scaled so that $\rho(A^{n}B)=1$, then $BA^{n}$ has an eigenvector $\tilde{v}_{0}=BA^{n}\tilde{v}_{0}$ close to $v_{0}$. For $i\geq 0$, let $\tilde{v}_{i}\coloneq A^{i}\tilde{v}_{0}$. For each $i$, $\tilde{v}_{i}$ must be close to $v_{i}$. So, for sufficiently small perturbation, the vectors ±$\pm\tilde{v}_{0},\dots,\pm\tilde{v}_{n}$ are the vertices of a convex polygon $\tilde{S}$. Note that the combinatorics (6.2) are unchanged, so $\tilde{S}$ is an invariant polygon for $(A,B)$ and the previous arguments persist. Therefore $A^{n}B$ is still the unique SMP. ∎
7. What are the SMPs if $(A,B)\in\mathcal{R}_{neg}$?
In this section we will assume $A,B$ both have negative determinants. Notice that this immediately implies both $A,B$ are real diagonalizable.
Lemma 7.1.
If $A,B$ are matrices with negative determinants and $AB$ is not real diagonalizable, then $(A,B)$ is crossing or reducible.
Proof.
To simplify notation, we will be using the associated $5$-tuple as defined in (2.1). If $AB$ is not real diagonalizable then $z^{2}-4uv\leq 0\implies-2\sqrt{uv}\leq z\leq 2\sqrt{uv}$. However, this implies
$4uv-uy^{2}-vx^{2}+xyz-z^{2}\geq-uy^{2}-vx^{2}-2xy\sqrt{uv}=(\sqrt{-u}y-\sqrt{-v}x)^{2}\geq 0$
Hence
$4uv-uy^{2}-vx^{2}+xyz-z^{2}\geq 0$
Which means $(A,B)$ is crossing or reducible by combining Theorems 2.1, 2.2 and 4.1. ∎
We are now ready to prove the 3rd part of Theorem 1.9.
Theorem 7.2.
If both $A,B$ have negative determinants, are irreducible and $A^{2}\not=\operatorname{JSR}(A,B)^{2}I$, $B^{2}\not=\operatorname{JSR}(A,B)^{2}I$, then all the SMPs are among $A,B,AB$.
Proof.
Without loss of generality we will normalize both $A,B$ so that $\operatorname{JSR}(A,B)=1$. If $(A,B)$ is crossing we are done by Corollary 5.4, thus we may assume $(A,B)$ is not crossing. By Lemma 7.1, $AB$ must be real diagonalizable. Now by applying Lemma 6.7 twice (Interchanging the role of $A$ and $B$), we see that both $(A,AB)$ and $(B,AB)$ are crossing. However, by applying Lemma 6.8 to the pair $(A,AB)$ and $(B,AB)$ we see that all SMPs must be of the form $AB^{n}$ or $B$ as well as of the form $A^{n}B$ or $A$, unless $A$ or $B$ have finite order (Since $\operatorname{JSR}(A,B)=1$). However, the only matrices that have finite order and negative determinant must be conjugate to reflection and so have order $2$, which by assumption is not the case. Thus all the SMPs must be either $A,B,AB$ as those are the only products the two lists above have in common. ∎
8. Ergodic Optimization and Joint Spectral Radius
To tackle the co-parallel case we need to view the joint spectral radius as a quantity related to ergodic optimization. Given a finite set of matrices $\mathcal{A}=\{A_{0},\dots,A_{k-1}\}$, let $\sigma$ be the shift on $\Sigma_{k}=\{0,\dots,k-1\}^{\mathbb{N}}$. We define the following map for $\omega=(\omega_{i})_{i\in\mathbb{N}}$
$\mathcal{L}(\omega,n)=A_{\omega_{n}}\cdots A_{\omega_{1}}$ (8.1)
for which we may define the Lyapunov exponent.
Definition 8.1.
Let $\mathcal{M}_{\sigma}$ be the set of shift invariant probability measures. For each $\mu\in\mathcal{M}_{\sigma}$, the Lyapunov exponent of $\mathcal{L}$ with respect to $\mu$ is defined as
$\lambda(\mu)\coloneq\lim_{n\to\infty}\frac{1}{n}\int_{\Sigma_{k}}\log\|\mathcal{L}(\omega,n)\|d\mu(\omega).$ (8.2)
The joint spectral radius and the Lyapunov exponent of $\mathcal{L}$ are connected by the following abbreviated version of a theorem by Morris:
Theorem 8.2.
[21]
$\log\operatorname{JSR}(\mathcal{A})=\sup_{\mu\in\mathcal{M_{\sigma}}}\lambda(\mu)$ (8.3)
The supremum is always attained for some ergodic measure.
Any measure $\mu$ attaining the supremum is called a maximizing measure.
Note 8.3.
Each SMP gives rise to a maximizing measure supported on a periodic orbit. Two SMPs give rise to the same maximizing measure if and only if one is a cyclic permutation of the other.
To define Sturmian maximizing measures, we first define Sturmian sequences:
Definition 8.4.
Given $\gamma,\rho\in[0,1]$ we define two sequences in $\Sigma_{2}$:
$(\omega_{\gamma,\rho}^{-})_{n}$ $=\lfloor\gamma(n+1)+\rho\rfloor-\lfloor\gamma n+\rho\rfloor$
$(\omega_{\gamma,\rho}^{+})_{n}$ $=\lceil\gamma(n+1)+\rho\rceil-\lceil\gamma n+\rho\rceil$
The sequence $\omega_{\gamma,\rho}^{-}$ is called lower Sturmian and $\omega_{\gamma,\rho}^{+}$ is called upper Sturmian of slope $\gamma$ and intercept $\rho$.
Note 8.5.
These are also called mechanical words. See [17, p.53] for a more details.
The theorem below is, to our knowledge, folklore and describes all the properties of Sturmian sequences and measures we will need. Most of this theorem is proved in detail in [6].
Theorem 8.6 (Folklore).
Denote $X_{\gamma}$ as the set of all Sturmian sequences of slope $\gamma$. Then we have the following properties:
  1. (1)
    $X_{\gamma}$ is a compact and invariant under the shift.
  2. (2)
    The shift is uniquely ergodic on $X_{\gamma}$. This measure $\mu_{\gamma}$ is called the Sturmian measure of parameter $\gamma$.
  3. (3)
    The map $\gamma\mapsto\mu_{\gamma}$ where $\gamma\in[0,1]$ is continuous. Hence the set of all Sturmian measures is compact.
9. What are the possible SMPs if $(A,B)\in\mathcal{R}_{copar}$
In this section, we describe the possible SMPs given that $(A,B)$ is co-parallel. We remind the reader, that given a word $w\in\{0,1\}^{*}$ and matrices $A,B$, we denote $w(A,B)$ as the matrix product given by replacing each 00 with $A$ and $1$ with $B$. We begin this section with introducing special words that come from Sturmian sequences.
Definition 9.1.
Let $\frac{p}{q}\in\mathbb{Q}\cap[0,1]$ with $\gcd(p,q)=1$ and $\rho\in[0,1]$.
  • $q$-prefix of $\omega_{\frac{p}{q},\rho}^{-}$ is called a Sturmian word
  • $q$-prefix of $\omega_{\frac{p}{q},0}^{-}$ is called a Christoffel word
Note 9.2.
All Sturmian words are primitive. Christoffel words are exactly the Sturmian words that are Lyndon. When we refer to a Sturmian/Christoffel product we mean a product $w(A,B)$ where $w$ is a Sturmian/Christoffel word.
All Christoffel words can be iteratively obtained using the Christoffel tree which is constructed as follows. Each node will be a pair of words. The root is $(0,1)$. Each node $(u,v)$ has two children, the left child is $(u,uv)$ and the right child is $(uv,v)$( See Fig. 3 for the first three layers of the tree.).
Theorem 9.3.
[2] The Christoffel tree contains exactly once the decomposition $(u,v)$ of each Christoffel word $w=uv$ of length greater than $1$. All words in each node are Christoffel.
[…]
[.$(0,1)$ [.$(0,01)$ [.$(0,001)$ ] [.$(001,01)$ ] ] [.$(01,1)$ [.$(01,011)$ ] [.$(011,1)$ ] ] ]
Figure 3. The Christoffel tree
We now prove a few lemmas that will be needed in the main theorem.
Lemma 9.4.
If $(A,B)$ is co-parallel, then $AB$ is co-parallel to both $A$ and $B$, with its axis lying between the axis of $A$ and of $B$. Furthermore $(A^{n},B^{k})$ is also co-parallel for any $n,k\geq 1$ with the same axis as $A$ and $B$.
Proof.
The first part is proved in [14, Lemma 4.2]. The second part follows from the fact that $A^{n}$ and $B^{k}$ have exactly the same eigenvectors as $A$ and $B$. The only way that would not be the case is if $A^{n}$ or $B^{k}$ are multiples of identity. However, that is impossible since that would mean $A$ or $B$ are conjugate to a multiple of a rational rotation and therefore not real diagonalizable. ∎
To make the notation more manageable in the following corollary $u$ should be understood as the product $u(A,B)$.
Corollary 9.5.
Let $(A,B)$ be co-parallel. Then a Christoffel product $w(A,B)$ is co-parallel to any other Christoffel product $w^{\prime}(A,B)$.
Proof.
We show any product the Christoffel tree is co-parallel to any other product in the tree. Using the Lemma 9.4, it follows by induction that given any node $(u,v)$, for all of its descendants $(u^{\prime},v^{\prime})$, $u^{\prime}$ (resp. $v^{\prime}$) is co-parallel with both $u$ and $v$ with its axis laying between the axis of $u$ and $v$.
Now take any two distinct nodes $(u_{1},v_{1})$ and $(u_{2},v_{2})$ (We assume one is not a descendant of another, similar argument works here as well) and trace them back to the last common ancestor say $(u^{\prime},v^{\prime})$. Without loss of generality, $(u_{1},v_{1})$ is a descendant of $(u^{\prime},u^{\prime}v^{\prime})$ and $(u_{2},v_{2})$ is a descendant $(u^{\prime}v^{\prime},v^{\prime})$, otherwise there would be a deeper node in common. However, the axis of $u_{1}v_{1}$ must be between both axis of $u^{\prime},u^{\prime}v^{\prime}$ and axis of $u_{2}v_{2}$ must be between both axis of $u^{\prime}v^{\prime},v^{\prime}$. Therefore, they must be disjoint. Hence $u_{1}v_{1}$ and $u_{2}v_{2}$ must be co-parallel. ∎
Lemma 9.6.
Let $(A_{0},A_{1})$ be co-parallel. Then the associated Lyapunov exponent $\lambda:\mathcal{M}_{\sigma}\to\mathbb{R}$, as defined in Definition 8.1, is continuous.
Proof.
We may assume $A_{0},A_{1}$ have non-negative trace by possibly negating the matrices. Since they have non-negative trace and positive determinants, they have positive eigenvalues. Hence, there exists a strictly forward invariant cone $C$ such that $A_{0}(C)\subset C$ and $A_{1}(C)\subset C$; any cone that contains the attracting eigenvectors in its interior and does not contain the repelling eigenvectors will work. Existence of such a cone is equivalent to the associated cocycle $\mathcal{L}(\omega,n)$ being dominated (see [3] for proof and definitions). Dominance guarantees existence of a continuous map $s:\Sigma_{2}\to\mathbb{R}^{2}$ such that $\|s(\omega)\|=1$ and:
$\frac{\mathcal{L}(\omega,n)s(\omega)}{\|\mathcal{L}(\omega,n)s(\omega)\|}=s(\sigma^{n}(\omega)).$
Furthermore, using [5, Corollary 2.4] we know that given $\mu\in\mathcal{M}_{\sigma}$, for $\mu$ almost every $\omega$:
$\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{L}(\omega,n)\|=\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{L}(\omega,n)s(\omega)\|$
Therefore, integrating and applying dominated convergence theorem:
$\lambda(\mu)=\int_{\Sigma_{2}}\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{L}(\omega,n)s(\omega)\|d\mu(\omega).$
Using the properties of $s(\omega)$ the limit can be re-written:
$\lim_{n\to\infty}\frac{1}{n}\log\|\mathcal{L}(\omega,n)s(\omega)\|=\lim_{n\to\infty}\frac{1}{n}\sum_{i=0}^{n-1}\log\|\mathcal{L}(\sigma^{j}(\omega),1)s(\sigma^{j}(\omega))\|$
Now, by Birkhoff ergodic theorem:
$\lambda(\mu)=\int_{\Sigma_{2}}\log{\|\mathcal{L}(\omega,1)s(\omega)\|}d\mu(\omega)$
Since $\log{\|\mathcal{L}(\omega,1)s(\omega)\|}$ is continuous, $\lambda$ is continuous with respect to the weak* topology. ∎
We now state another theorem of Jorgensen-Smith, we remind the reader that $W(a,b)$ denotes the set of products that have $a$ $A$’s and $b$ $B$’s.
Theorem 9.7.
[14] If $(A,B)$ is co-parallel, then for any pair of positive integers $a$ and $b$, the products that maximize the spectral radius in $W(a,b)$ are cyclic permutations of the prefix of length $a+b$ of the product $\omega_{\frac{b}{a+b},0}^{-}(A,B)$.
We note that Jorgensen and Smith proved this result under the additional hypothesis $(A,B)\in SL(2,\mathbb{R})^{2}$ and $\operatorname{tr}{A},\operatorname{tr}{B}>2$. However, the more general statement follows by re-scaling the matrices and noting that all products in $W(a,b)$ have the same determinant. Also, Jorgensen and Smith maximize the trace, but under the hypothesis this is equivalent to maximizing the spectral radius.
We are finally ready to prove 4th part of Theorem 1.9.
Theorem 9.8.
Let $(A,B)$ be co-parallel. There always exists exactly one Sturmian maximizing measure. If an SMP exists, it must be unique (up to cyclic permutations) and be Sturmian.
Proof.
We first show that there is always exactly one Sturmian maximizing measure. By Theorem 1.2 we know that the maximizing measures can be approximated arbitrarily well by periodically supported measures. In this context, Theorem 9.7 says that the Lyapunov exponent over periodic measures of a given period is maximized for some Sturmian measure of a rational parameter. Denote $\mathcal{M}_{\sigma}$ as all shift invariant probability measures, $\mathcal{M}^{p}_{\sigma}$ as all shift invariant probability measures supported on a periodic orbit and $\mathcal{M}^{\mathbb{Q}}_{\sigma}$ as all Sturmian measures of rational parameter. Combining the two results we get the following chain of equalities.
[…]
However, by Theorem 8.6 the set of all Sturmian measures is compact. Therefore we know there is a convergent sequence of Sturmian measures of rational parameter $\mu_{n}\to\mu$ that converges to a maximizing Sturmian measure.
To show uniqueness, we prove that the map $f(t)=\lambda(\mu_{t})$ is strictly concave on $[0,1]$, where $\mu_{t}$ denote the Sturmian measure of parameter $t$. Given any two rational numbers $t_{1},t_{2}\in\mathbb{Q}$, they can be written with a common denominator, $t_{1}=\frac{p_{1}}{q},t_{2}=\frac{p_{2}}{q}$. Let $w_{1}$ be the prefix of length $q$ of the Sturmian sequence $\omega_{t_{1},0}^{-}$ and $w_{2}$ be the prefix of length $q$ of the Sturmian sequence $\omega_{t_{2},0}^{-}$. Lastly, we let $w_{3}$ be the prefix of length $2q$ of the Sturmian sequence $\omega_{\frac{t_{1}+t_{2}}{2},0}^{-}$. Notice that $w_{1},w_{2},w_{3}$ are all powers of Christoffel words. Denote $\Pi_{1}=w_{1}(A,B),\Pi_{2}=w_{2}(A,B),\Pi_{3}=w_{3}(A,B)$. By Corollaries 9.5 and 9.4 $(\Pi_{1},\Pi_{2})$ is co-parallel. Applying Corollary 4.5:
$\rho(\Pi_{1}\Pi_{2})>\rho(\Pi_{1})\rho(\Pi_{2})$
Furthermore, by Theorem 9.7:
$\rho(\Pi_{3})\geq\rho(\Pi_{1}\Pi_{2})$
Combining the two inequalities we get
$\rho(\Pi_{3})>\rho(\Pi_{1})\rho(\Pi_{2})$
Taking both sides to the power $\frac{1}{2q}$ and taking the logarithm we get:
$\log(\rho(\Pi_{3})^{\frac{1}{2q}})>\frac{\log(\rho(\Pi_{1})^{\frac{1}{q}})}{2}+\frac{\log(\rho(\Pi_{2})^{\frac{1}{q}})}{2}$
The above inequality can be re-written in terms of $f$:
$f\left(\frac{p_{1}+p_{2}}{2q}\right)>\frac{f\left(\frac{p_{1}}{q}\right)+f\left(\frac{p_{2}}{q}\right)}{2}.$
Which shows that $f$ is strictly mid-point concave on the rationals. Since $f$ is continuous by Lemma 9.6, it must be strictly mid-point concave everywhere. Hence, $f$ is strictly concave. Any continuous strictly concave function on a compact set has exactly one maximum. This proves that there is exactly one Sturmian maximizing measure.
We are now ready to show that if an SMP exists, then it is Sturmian and is unique up to cyclic permutations. We know by Theorem 9.7 that all the SMPs must be Sturmian since they have to maximize spectral radius in their respective set $W(a,b)$. Therefore, they all must come from some Sturmian maximizing measure of rational parameter. However, these are unique, so there can only be one such SMP up to cyclic permutations. ∎
Note 9.9.
We suspect that the maximizing measure is always unique (and therefore Sturmian) in the co-parallel region. We also suspect that the counterexamples to the finiteness conjecture form a subset of zero Hausdorff dimension. Some results in this direction were given by [6, 11, 13].
10. Generic Uniqueness
In this section, we prove that in the union of the three regions $\mathcal{R}_{cross}\cup\mathcal{R}_{mix}\cup\mathcal{R}_{neg}$, the SMPs are generically unique and of the form given in Theorem 1.9. We first prove a few lemmas. We denote $\lambda_{n}$ as the $n$-dimensional Lebesgue measure.
Lemma 10.1.
Consider a polynomial map $\pi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ such that $\lambda_{m}(\pi(\mathbb{R}^{n}))>0$. Then for any measurable set $E\subseteq\mathbb{R}^{m}$,
$\lambda_{m}(E)=0\implies\lambda_{n}(\pi^{-1}(E))=0$
Proof.
Denote $C_{\pi}\subseteq\mathbb{R}^{n}$ as the critical points of $\pi$. Since the map $\pi$ is a polynomial, $C_{\pi}$ is algebraic and so either $\lambda_{n}(C_{\pi})=0$ or $C_{\pi}=\mathbb{R}^{n}$. By Sard’s theorem $\lambda_{m}(\pi(C_{\pi}))=0$. Since $\lambda_{m}(\pi(\mathbb{R}^{n}))>0$, $C_{\pi}$ must have zero measure. This means $\pi$ is almost everywhere a local submersion and now the lemma follows by rank theorem and Fubini, for details see [24]. ∎
We apply the above lemma to the mapping:
$(A,B)\mapsto(\operatorname{tr}{A},\operatorname{tr}{B},\operatorname{tr}{AB},\det{A},\det{B}).$
By Theorem 2.3 its image has positive measure. Thus, if we want to prove a generic statement about pairs of real matrices $(A,B)$, we may equivalently prove it for generic $5$-tuples. We will tacitly do so from now on, without repeatedly invoking the above lemma.
Lemma 10.2.
For Lebesgue almost every pair $(A,B)\in\mathcal{R}_{cross}\cup\mathcal{R}_{mix}\cup\mathcal{R}_{neg}$, all the SMPs (up to cyclic permutation) are of the form given by Theorem 1.9.
Proof.
If $(A,B)\in\mathcal{R}_{cross}$, all the SMPs are either $A$ or $B$ by Corollary 5.4 so there is nothing left to prove.
If $(A,B)\in\mathcal{R}_{mix}$, we may assume that $A,B$ are invertible as non-invertible matrices form a null set. Now by Theorem 6.9 it is sufficient to show that pairs $(A,B)$ for which there exists $k\geq 1$ such that
$A^{k}=\operatorname{JSR}(A,B)^{k}I,$
form a measure zero set. Assume $(A,B)$ are such that the above holds, then $A$ must be conjugate to a multiple of a rotation by $2\pi\theta$ where $\theta\in\mathbb{Q}$. Such matrices satisfy the equation $\det(A)=\frac{(\operatorname{tr}{A})^{2}}{4\cos^{2}(2\pi\theta)}$. Thus, for fixed $\theta$, all associated $5$-tuples live on the on the parametric family
$M_{\theta}=(x,y,z,\frac{x^{2}}{4\cos^{2}(2\pi\theta)},v)$
which has Lebesgue zero measure in $\mathbb{R}^{5}$. Since the set of rational numbers is countable the set $\cup_{\theta\in\mathbb{Q}}M_{\theta}$ also has measure zero. Thus, such $(A,B)$ must form a measure zero set.
If $(A,B)\in\mathcal{R}_{neg}$, then by Theorem 7.2 it is sufficient to show that pairs $(A,B)$ such that
$A^{2}=\operatorname{JSR}(A,B)^{2}I\quad\text{or }\quad B^{2}=\operatorname{JSR}(A,B)^{2}I$
form a measure zero set. Without loss of generality assume $A$ satisfies the above, then $A$ must be conjugate to a multiple of a reflection. Such matrices satisfy the equation $\operatorname{tr}{A}=0$. Hence, all associated $5$-tuples live in
$(0,y,z,u,v)$
which is measure zero. Similarly if $B$ is conjugate to reflection we get a zero measure set. Hence, such $(A,B)$ must form a measure zero set.
Lemma 10.3.
The set of reducible pairs of matrices $(A,B)$ has measure zero.
Proof.
By Theorem 2.1, the $5$-tuples of reducible pairs of matrices is the pre-image of $\{0\}$ of a particular non-zero polynomial. Since the zero set of any non-zero polynomial must has have zero measure, the lemma follows. ∎
Lemma 10.4.
For Lebesgue almost every pair $(A,B)$ of real $2\times 2$ matrices, if $\Pi$ is an SMP that has non-real eigenvalues, then $\Pi=A$ or $\Pi=B$ and the SMP is unique.
Proof.
By Lemma 10.3 we may assume $(A,B)$ is irreducible. For those pairs, $\varrho\coloneqq\mathrm{JSR}(A,B)$ is nonzero and there exists a centrally symmetric convex body $C\subseteq\mathbb{R}^{2}$ such that the convex hull of $A(C)\cup B(C)$ is $\varrho C$ (See [25]). Suppose $\Pi$ is an SMP with non-real eigenvalues. Then, $\det\Pi=\varrho^{2k}$, where $k\coloneqq|\Pi|$, so the inclusion $\Pi(C)\subseteq\varrho^{k}C$ is actually an equality. It follows that either $A(C)$ or $B(C)$ equals $\varrho C$. Almost surely we have $\det A\neq\det B$, and in this case $A(C)\neq B(C)$. Since $\Pi$ is primitive, it must be either $A$ or $B$, and $\Pi$ is the unique SMP. ∎
Given a word $w\in\{0,1\}^{*}$, define its signature $(m,k,l)$ as follows: we first cyclically permute $w$ so that it is a Lyndon word and the three numbers are: $m$ is the number of 00’s, $k$ is the number of $1$’s, and $l$ is the number of $01$’s that can be found within that new word. We remind the reader that given a word $w$,$F_{w}(x,y,z,u,v)$ denotes the Fricke polynomial described in Theorem 2.5.
Lemma 10.5.
Let $w$ and $w^{\prime}$ be words whose signatures $(a,b,c)$ and $(a^{\prime},b^{\prime},c^{\prime})$ are linearly independent, then the polynomials $F_{w}(x,y,z,0,0),F_{w^{\prime}}(x,y,z,0,0)$ are algebraically independent. That is, there is no non-zero polynomial $P$ such that:
$P(F_{w}(x,y,z,0,0),F_{w^{\prime}}(x,y,z,0,0))=0$
Proof.
Consider a pair of matrices $A,B$ associated to the $5$-tuple $(x,y,z,0,0)$ where $x,y,z$ are free. Now we use observations similar ones in Lemma 6.6. Cyclically permuting $w$ if necessary, $w(A,B)$ can be written as
$w(A,B)=A^{a_{1}}B^{b_{1}}\cdots A^{a_{c}}B^{b_{c}}$
with $a_{i},b_{i}\geq 1$. Since $A,B,AB$ are all non-invertible, by Cayley-Hamilton we have
$A^{n}=(\operatorname{tr}{A})^{n-1}A,\quad B^{n}=(\operatorname{tr}{B})^{n-1}B,\quad(AB)^{n}=(\operatorname{tr}{AB})^{n-1}AB.$
Thus, we may reduce all powers of $A$ and $B$ by pulling out traces:
$w(A,B)=(\operatorname{tr}{A})^{a-c}(\operatorname{tr}{B})^{b-c}(AB)^{c}$
We now take the trace and get the following expression:
$\operatorname{tr}{w(A,B)}=(\operatorname{tr}{A})^{a-c}(\operatorname{tr}{B})^{b-c}(\operatorname{tr}{AB})^{c}=x^{a-c}y^{b-c}z^{c}$
By exactly the same reasoning:
$\operatorname{tr}{w^{\prime}(A,B)}=(\operatorname{tr}{A})^{a^{\prime}-c^{\prime}}(\operatorname{tr}{B})^{b^{\prime}-c^{\prime}}(\operatorname{tr}{AB})^{c^{\prime}}=x^{a^{\prime}-c^{\prime}}y^{b^{\prime}-c^{\prime}}z^{c^{\prime}}$
Now if the polynomials $F_{w}(x,y,z,0,0),F_{w^{\prime}}(x,y,z,0,0)$ were algebraically dependent then:
$F_{w}(x,y,z,0,0)=x^{a-c}y^{b-c}z^{c}\text{ and }F_{w^{\prime}}(x,y,z,0,0)=x^{a^{\prime}-c^{\prime}}y^{b^{\prime}-c^{\prime}}z^{c^{\prime}}$
would also have to be dependent. However, since $(a,b,c)$ and $(a^{\prime},b^{\prime},c^{\prime})$ are linearly independent, we arrive at a contradiction. ∎
We are now ready to prove generic uniqueness.
Theorem 10.6.
Almost every pair $(A,B)\in\mathcal{R}_{cross}\cup\mathcal{R}_{mix}\cup\mathcal{R}_{neg}$ has a unique SMP that is of the form described in Theorem 1.9.
Proof.
Assume for a contradiction that there exists a subset $\mathcal{P}$ of $\mathcal{R}_{cross}\cup\mathcal{R}_{mix}\cup\mathcal{R}_{neg}$ with positive Lebesgue measure such that every pair in $\mathcal{P}$ admits two SMPs. We know for almost every pair $(A,B)$ all SMPs are of the form $A^{n}B$ or $AB^{n}$ by Lemma 10.2. Since the set of such products is countable, we can find a pair of distinct primitive words $w$, $w^{\prime}$ of the form $10^{n}$ 1 or $01^{n}$ and a subset $\mathcal{Q}\subseteq\mathcal{P}$ of positive measure such that for every $(A,B)\in\mathcal{Q}$, $w(A,B)$ and $w^{\prime}(A,B)$ are both SMPs. By Lemma 10.4 we may reduce the set $\mathcal{Q}$ if necessary, and assume $w(A,B),w^{\prime}(A,B)$ have real eigenvalues for all $(A,B)\in\mathcal{Q}$.
Fix $(A,B)\in\mathcal{Q}$ and consider the two distinct SMPs $\Pi=w(A,B)$ $\Pi^{\prime}=w^{\prime}(A,B)$ with lengths $k\coloneqq|\Pi|$, $k^{\prime}\coloneqq|\Pi^{\prime}|$. Let $\lambda_{1}$, $\lambda_{2}$ (resp. $\lambda_{1}^{\prime}$, $\lambda_{2}^{\prime}$) be the eigenvalues of $w$ (resp. $w^{\prime}$). By assumption, all eigenvalues $\lambda_{1},\lambda_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$ are real. Furthermore,
$\mathrm{JSR}(A,B)=\max\left\{|\lambda_{1}|^{\frac{1}{k}},|\lambda_{2}|^{\frac{1}{k}}\right\}=\max\left\{|\lambda_{1}^{\prime}|^{\frac{1}{k^{\prime}}},|\lambda_{2}^{\prime}|^{\frac{1}{k^{\prime}}}\right\}\,.$ (10.1)
Since eigenvalues are real,
$\big{(}\lambda_{1}^{2k^{\prime}}-(\lambda^{\prime}_{1})^{2k}\big{)}\big{(}\lambda_{1}^{2k^{\prime}}-(\lambda^{\prime}_{2})^{2k}\big{)}\big{(}\lambda_{2}^{2k^{\prime}}-(\lambda^{\prime}_{1})^{2k}\big{)}\big{(}\lambda_{2}^{2k^{\prime}}-(\lambda^{\prime}_{2})^{2k}\big{)}=0\,.$ (10.2)
Expanding and using Newton–Girard identities, this equation becomes:
$P(\operatorname{tr}\Pi,\operatorname{tr}\Pi^{\prime},\det\Pi,\det\Pi^{\prime})=0\,,$ (10.3)
where $P$ is some polynomial in four variables, with real coefficients. Letting as usual $x=\operatorname{tr}A$, $y=\operatorname{tr}B$, $z=\operatorname{tr}AB$, $u=\det A$, $v=\det B$, the relation above can be rewritten
$P\big{(}F_{w}(x,y,z,u,v),F_{w^{\prime}}(x,y,z,u,v),u^{a}v^{b},u^{a^{\prime}}v^{b^{\prime}})=0$ (10.4)
where $F_{w},F_{w^{\prime}}$ are the Fricke polynomials and $a,b,a^{\prime},b^{\prime}$ are appropriate non-negative integers.
Equation (10.4) holds for all $(A,B)$ in a positive measure subset of pairs $(A,B)$. It follows that the same equation, seen in terms of variables $x,y,z,u,v$, holds on a positive measure subset of $\mathbb{R}^{5}$. Since proper algebraic sets have zero measure, we conclude that the equation holds on all of $\mathbb{R}^{5}$. Setting $u=v=0$ we obtain:
$P\big{(}F_{w}(x,y,z,0,0),F_{w^{\prime}}(x,y,z,0,0),0,0)=0.$
However, since $w,w^{\prime}$ are two distinct words of forms $10^{n}$ 1 or $10^{n}$, they necessarily have linearly independent signatures. By Lemma 10.5 such $P$ cannot exits and we arrive at a contradiction. ∎
With this theorem, we have completed the proof of Theorem 1.9.
11. The unexplored regions
There are essentially two regions that remain unexplored $\mathcal{S}_{anti},\mathcal{S}_{complex}$:
  • $\mathcal{S}_{anti}$ is the region where $(A,B)$ is anti-parallel.
  • $\mathcal{S}_{complex}$ is the region where $A$ or $B$ have complex eigenvalues.
Some notes about these regions:
  • Generic uniqueness of SMPs (modulo only cyclic permutations) does not hold in $\mathcal{S}_{complex}$, see [4]. There are two cases, either both or only one matrix has complex eigenvalues. The examples in [4] are all in the latter category; however there exist examples where both matrices have complex eigenvalues. One such example is given by the $5$-tuple:
    $(-3.76601,-0.49459,-8.13153,3.52510,8.71249).$
  • By our classification, generically, non-Sturmian SMPs can only occur in these unexplored regions. It is not too hard to find explicit examples of non-Sturmian SMPs in both of them.
[…]
12. Non-negative matrices
In the introduction, we claimed that our classification encompasses all non-negative matrices. We are now ready to prove this result. We begin with a lemma characterizing non-negative matrices with positive determinants.
Lemma 12.1.
Let $A,B\in GL^{+}(2,\mathbb{R})$ be irreducible real diagonalizable, then $(A,B)$ is simultaneously conjugate to non-negative matrices if and only the $(A,B)$ is either crossing or co-parallel and $\operatorname{tr}(A),\operatorname{tr}(B)\geq 0$
Proof.
We first prove the forward direction. We may assume our matrices $A,B$ are non-negative. By Perron-Frobenius theorem, both Möbius transformations $f_{A}$ and $f_{B}$ have an attracting non-negative fixed point and a repelling non-positive fixed point. Hence $(A,B)$ cannot be anti-parallel. Since $A,B$ are non-negative, it is clear that $\operatorname{tr}{A},\operatorname{tr}{B}\geq 0$ and so this direction follows.
Now assume that $(A,B)$ is co-parallel or crossing with non-negative traces. Since the traces are non-negative and determinants are positive, both $A$ and $B$ have only positive eigenvalues. Therefore, in either case (co-parallel or crossing) there exists a cone $C$ such that $A(C)\subseteq C$ and $B(C)\subseteq C$ (Any cone that contains the dominant eigenvectors of $A$ and $B$ and does not contain the non-dominant eigendirections of $A$ and $B$ will work). Now take a matrix $P$ that maps the cone $C$ to the first quadrant and consider $(PAP^{-1},PBP^{-1})$. These new matrices keep the first quadrant invariant by construction. That means both of them must be non-negative as desired. ∎
We are now ready to prove the following corollary:
Corollary 12.2.
Let $(A,B)$ be matrices that are simultaneously conjugate to non-negative matrices, then their SMP, if exists, must be Sturmian.
Proof.
By Theorem 1.9 if either $A$ or $B$ have a negative determinant, then an SMP exists and must be Sturmian since products of the form $A^{n}B$ and $AB^{n}$ are Sturmian. If both $A,B$ have positive determinants then by Lemma 12.1 $(A,B)$ must be crossing or co-parallel. However, again invoking Theorem 1.9 we see that if an SMP exists, it must be Sturmian. ∎
Acknowledgements
I would like to thank my advisor Jairo Bochi for invaluable suggestions and comments. This paper would not be possible without his insight and encouragement.